# How do Laws Produce the Future?

### v1

@@ 6 Electromagnetism, Second Pass
The wave equation for the scalar potential (17) fits the form of a three-dimensional wave equation (30) and thus we can formulate it as a spatiotemporally local law like (40),
@@
The wave equation for the vector potential (18) is three separate three-dimensional wave equations of the earlier form (30), one for each component of the potential. We can formulate it as
@@
For point charges, current density $\vec{J}$ that sources the vector potential can be formed by combining contributions of the form $\rho\vec{v}$ associated with each point charge, where $\rho$ is a delta function centered on the charge’s location and $\vec{v}$ is the charge’s velocity (Landau & Lifshitz,, 1971, sec. 28). As before, the velocity can be understood as a past derivative of the body’s location and thus depends on an arbitrarily short segment of its past trajectory. If the body has never moved faster than the speed of light, then this trajectory will lie within the past light-cone and the velocity will be a past light-cone neighborhood property. The current density will thus also be a past light-cone neighborhood property. What if charged matter is modeled instead by a charge distribution? In that case, the current density at a moment cannot be read off from the present charge distribution and its recent history. Maudlin, (2018, pg. 8) gives the example of a uniform sphere that maintains a uniform charge density. This is compatible with there being no current density or with their being a current density corresponding to a rotation of the charge about some axis. Thus, unlike the velocity of a point charge, the current density of a continuum cannot be reduced to the historic behavior of charge. One option would be to treat the charge distribution and current density as distinct properties of the instantaneous state. The current density then becomes a space-time point property, not a past light-cone neighborhood property. Again, (43) counts as a spatiotemporally local dynamical law.
In this paper, we have assumed that velocity should be reduced to changes in position over time. There is an alternative view available according to which instantaneous velocity is a genuine feature of the world at a moment that ought not be reduced to changes in position. Lange, (2005) calls this view “velocity primitivism.” As he puts it, “instantaneous velocity and trajectory are related only by virtue of natural law, not by metaphysical necessity.” That is, $\vec{v}=\frac{d\vec{x}}{dt}$ becomes a law of nature and not a definition. For the current density of a continuum, we have ended up in a similar position. Current density and charge density will be related by a law of nature: the continuity equation for charge,
$\frac{\partial\rho}{\partial t}=-\vec{\nabla}\cdot\vec{J}\ .$ (44)
This equation can be derived from Maxwell’s equations and should be derivable from the fundamental laws governing the evolution of the charged matter. [^34] For example, if matter is modeled as a classical Klein-Gordon or Dirac field then conservation of charge can be derived from the wave equations governing these fields. Thus, the continuity equation for charge might be regarded as a non-fundamental law (that, because of the divergence, is not spatiotemporally local—which is fine for a non-fundamental law).
@@
In words: the force from the electromagnetic field in a given direction depends on (i) the way the scalar potential is changing forward and backward in that direction along the past light-cone, (ii) the dot product of the charge’s velocity with the way the vector potential is changing forward and backward in that direction along the past light-cone, and (iii) the time derivative of the vector potential along the past trajectory of the charge.
@@
From (50) and (51), you can see that no matter how much force you exert on a massive body to increase its momentum, you cannot accelerate it beyond the speed of light—the momentum approaches infinity as the speed approaches $c$. Because the particles do not move faster than $c$, the momentum $\vec{p}^{\,p}$ will be a past light-cone neighborhood property and the right-hand side of (50), $\left(\frac{d}{dt}\right)^{f}\vec{p}^{\,p}$, will be a past-to-future light-cone neighborhood property. If we assume that we are dealing with purely electromagnetic forces then, as we have already discussed, the Lorentz force (45) that would appear on the right-hand side of (50) is a past light-cone neighborhood property. Thus, the reaction law (50) has the correct form to be a spatiotemporally local dynamical law. As in section 3, you could combine the non-dynamical Lorentz force law (48) and the dynamical reaction law (50) into a single dynamical law (assuming that the only forces at play are electromagnetic forces). Making that move, the laws of electromagnetism would all be dynamical laws and there would be no need for any non-dynamical laws.
@@
If the matter is modeled using continuous distributions of charge instead of point charges, the reaction of matter to forces is more complicated and will depend on the nature of the matter. You would need to couple some theory of the matter with the above laws of electromagnetism to get a complete theory that could fit the relativistic time evolution paradigm. A neatly relativistic way to do this would be to model the matter using a classical Klein-Gordon or Dirac field in a Maxwell-Klein-Gordon or Maxwell-Dirac classical field theory. Then, one would need to show that the wave equations for these fields can be formulated in ways that fit the relativistic time evolution paradigm, as we have seen that the wave equations of electromagnetism can—(42) and (43). For the Klein-Gordon field, that is straightforward as we can simply carry over the techniques that were used for the electromagnetic field. [^36] You can prove the same kind of causality theorem for the Klein-Gordon equation as the one we discussed for the wave equation near the beginning of this section (Wald,, 1984, sec. 10.1; Strauss,, 2008, pg. 234, problem 8). For the Dirac field, it is not so clear how to formulate the wave equation as a spatiotemporally local dynamical law. Note that with either the Klein-Gordon or Dirac equation in place, the Lorentz force law (49) is not needed as a fundamental non-dynamical law (though it could still be regarded as a non-fundamental law giving the density of force exerted by the electromagnetic field on the Klein-Gordon or Dirac field [^37] The idea that fields can experience forces is discussed in Sebens, (2018, 2021b).).
@@
Putting it all together, we have arrived at a way of formulating electromagnetism that seems like it could fit the relativistic time evolution paradigm. For point charges, we can take the laws to be given by the wave equations for the potentials, (42) and (43), the Lorentz force law, (48), and the relativistic version of Newton’s second law (50)—or, (48) and (50) could be combined. However, there is danger on the horizon for that kind of theory. We could run into problems of self-interaction like those faced by Newtonian gravity with point masses (discussed in section 3). Just as the gravitational field becomes infinite as you approach a point mass and ill-defined at its location, the electric field becomes infinite as you approach a point charge and ill-defined at its location. This causes trouble for the standard Lorentz force law (14). Does our new formulation of the force law (48) face the same problems? For a single point charge $q$ at rest, we can use the scalar potential $\phi=\frac{q}{r}$ and set the vector potential to zero. The force law (48) yields
@@
We have seen that this formulation of electromagnetism with point charges satisfies the first two conditions of the relativistic time evolution paradigm: the dynamical and non-dynamical laws take the correct form. We have not yet settled whether the formulation satisfies the third condition as we have not yet settled whether the theory is deterministic. To do so, we must figure out whether, in general, an arbitrarily thin law-abiding time slice yields a unique law-abiding future. That requires looking at the coupled equations for the field and particle dynamics—not merely finding the field produced by a specified source or the force from a specified field. [^40] See Frisch, (2004, sec. 2); Frisch, (2005, pg. 32–35). Doing this honestly would require a solution to the aforementioned problems of self-interaction. There is another complication as well. Hartenstein & Hubert, (2021) have shown that specifying law-abiding combinations of electromagnetic field and point charge states at an instant generically yields infinities or discontinuities that propagate along the future light-cone (called “shock fronts”) and lead to ill-defined dynamics. [^41] See also Lazarovici, (2018, sec. 8.1). These shock fronts occur because the arbitrarily chosen field around a point charge does not capture the way that the charge’s past motion would have acted as a source for the field. For example, at some moment you might have a particle with an initial velocity surrounded by the Coulomb field of a stationary point charge. That is not compatible with any reasonable past. Our formulation of determinism (from section 2) requires that we start with an (arbitrarily short) time interval where the laws are obeyed and that starting point should help in addressing these problems. We are assuming that the interaction between particle and field has been well-behaved in the immediate past of a moment and then asking whether the future evolution is well-behaved and unique.
@@
Shifting now to consider electromagnetism with a continuous charge distribution, the wave equations for the potentials, (42) and (43), would again be laws. In addition, we would need law(s) determining the evolution of the charged matter. Although the Lorentz force density law, (49), can be formulated as a spatiotemporally local dynamical law, it will not be a fundamental law for certain classical completions of electromagnetism. In particular, if we model the evolution of matter using the Klein-Gordon or Dirac equations in Maxwell-Klein-Gordon or Maxwell-Dirac field theory, it is not needed. As was discussed earlier, the Klein-Gordon equation can be formulated as a spatiotemporally local dynamical law and one might find a way to do so for the Dirac equation. The Maxwell-Klein-Gordon and Maxwell-Dirac field theories could then satisfy the first two conditions of the relativistic time evolution paradigm. They may satisfy the third condition as well—the dynamics appears to be deterministic and, although self-interaction is present, there is no infinite or ill-defined self-interaction. [^42] See Sebens, (2023).
In this section, we have seen that electromagnetism in the Lorenz gauge can be formulated in a way that satisfies the first two conditions of the relativistic time evolution paradigm, with the dynamical and non-dynamical laws taking an appropriate form. For point charges, the laws would be (42), (43), (48), and (50) (or (48) and (50) could be combined into a single dynamical law). These laws are serviceable, but they are not the most elegant and it would be valuable to search for a better spatiotemporally local formulation of electromagnetism.
The third condition of the relativistic time evolution paradigm is complicated by problems of self-interaction, but determinism can potentially be achieved if we reject point charges and instead treat the charged matter as a continuous charge distribution (as is done in Maxwell-Klein-Gordon or Maxwell-Dirac field theory).
@@ 7 The Path Ahead
This article opened with a quote from Maudlin presenting the idea of dynamic production, a quote that comes from the end of the final chapter of his book The Metaphysics Within Physics. Here is that quote again, this time with more context:
“The non-Humean [dynamic production] package is, I think, much closer to the intuitive picture of the world that we begin our investigations with. Certainly, the fundamental asymmetry in the passage of time is inherent in our basic initial conception of the world, and the fundamental status of the laws of physics is, I think, implicit in physical practice. Both of the strands of our initial picture of the world weave together in the notion of a productive explanation, or account, of the physical universe itself. The universe, as well as all the smaller parts of it, is made: it is an ongoing enterprise, generated from a beginning and guided towards its future by physical law. …I don’t think that scientific results have, as yet, impeached the basic non-Humean picture, and no philosophical arguments give us reason to displace it. The metaphysics within physics is, as of now, non-Humean, and we can do no better as philosophers than embrace it.” (Maudlin,, 2007, pg. 182–183)
@@
The appeal of the dynamic production account of laws is that it seems to be such a natural way of understanding much of physics. We have seen here that, even for the parts of physics that seem most conducive to an interpretation in terms of dynamic production (Newtonian gravity and electromagnetism), there are serious challenges that arise. We have also seen that these challenges can be addressed. Being careful about the inputs and outputs, we can understand how the laws produce the future and how that production might be incorporated into a relativistic picture of reality. Let us now turn briefly to two physical theories that are more hostile to an interpretation in terms of dynamic production: quantum field theory and general relativity. The purpose of this quick treatment is just to suggest that the dynamic production account has a shot at being extended to our most successful physical theories, and should not be viewed as roadblocked [^43] Worries about potential roadblocks are raised in Wharton, (2015); Adlam, (2022a); Chen & Goldstein, (2022); Lam & Wüthrich, (2023). in a way that would make the kind of project pursued here a futile exercise.
@@
For the non-relativistic quantum mechanics of a fixed number of particles, there exist a handful of “interpretations of quantum mechanics” that give explicit proposals about the laws and ontology—such as GRW theory, Bohmian mechanics, the many-worlds interpretation, and the many interacting worlds approach. These four options can all be fit straightforwardly into the time evolution paradigm, though there are other “retrocausal,” “all-at-once,” or “non-Markovian” options [^44] See Wharton, (2014); Adlam, (2018, 2022a, 2022b, 2023); Friederich & Evans, (2023); Barandes, (2023). that cannot. Focusing on the four aforementioned options, the dynamics will be either first-order or second-order depending on the interpretation. The many-worlds interpretation simply takes the first-order Schrödinger equation to give the dynamics of the wave function. This is naturally interpreted as taking the wave function at a moment as input (with no need to look to the arbitrarily-short past) and returning as output the way the wave function changes into the arbitrarily short future (Easwaran,, 2014, pg. 857). GRW theory introduces stochastic exceptions to the Schrödinger equation that can be understood as giving probabilities for different behavior in the arbitrarily short future (Sebens,, 2015a, sec. 2). Bohmian mechanics is often presented as adding a first-order guidance equation that is another law, beyond the Schrödinger equation, determining the velocities of particles from their locations and the wave function (velocities that could be interpreted as future neighborhood properties). However, a second-order guidance equation can be used instead provided we impose the first-order guidance equation as a restriction on initial conditions (Goldstein & Struyve,, 2015; Dewdney,, 2023). In the many interacting worlds approach, the dynamics is second-order as particles are reacting to forces by Newton’s second law and (at the fundamental level) there is no wave function evolving by the Schrödinger equation (Hall et al.,, 2014; Sebens,, 2015b). The fact that quantum theories normally posit first-order dynamics has been used by Builes & Impagnatiello, (forthcoming) to argue that we live in a world where the present moment is enough to produce the future (the arbitrarily-short past is not needed as input)—in their words, “our universe is Markovian.” I prefer to leave open whether the fundamental dynamics will turn out to be first-order or second-order.
@@
Deciding between these competing interpretations of quantum mechanics is difficult. The many-worlds interpretation has been touted as “the only game in town” because it is the only option that can be immediately extended to relativistic quantum field theory (Wallace,, 2012, pg. 35; Wallace,, 2023). That virtue could be challenged either because the many-worlds interpretation has other serious problems (Adlam,, 2023) or because it is not so easy to formulate quantum field theory in terms of a quantum state evolving by a Schrödinger equation (Sebens,, 2022a). [^45] Adlam, (2022a, sec. 2.1) takes the use of path integrals in quantum field theory to suggest that the theory might best be understood outside the time evolution paradigm. Although it may be possible to use path integrals to formulate the theory in a way that violates the time evolution paradigm, path integrals in quantum field theory can be viewed as tools for calculating state evolution. The problem for Schrödinger evolution is that it is unclear whether the quantum state is a particle wave function (in Fock space) or a field wave functional, though hopefully one of these proposals can be made to work. Although the other interpretations may not be as straightforward to extend to quantum field theory as the many-worlds interpretation, there exist a number of promising strategies for doing so. The attempts to extend GRW and Bohmian mechanics to quantum field theory include non-local interactions that are incompatible with the relativistic time evolution paradigm (Maudlin,, 2019, ch. 7). This is to be expected, as it is the lesson of EPR and Bell’s theorem (taken along with the relevant experimental tests and making certain natural assumptions) that any single-world quantum theory must include such non-local interactions (Maudlin,, 2014, 2019). One might react to this situation by abandoning the relativistic time evolution paradigm and instead simply seeking theories that fit the original time evolution paradigm, accepting that dynamic production occurs relative to some preferred foliation. Alternatively, one might work to show that a theory with multiple worlds (like the many-worlds interpretation or the many interacting worlds approach) can be cast in a form that fits the relativistic time evolution paradigm by putting the ontology in spacetime [^46] See Wallace & Timpson, (2010); Sebens, (2015b, sec. 11); Sebens, (2021a, sec. 4.4). and showing that the dynamical laws are spatiotemporally local. Of course, one could also react to this situation by abandoning dynamic production (Adlam,, 2023). I simply want to highlight that there are ways to potentially retain dynamic production in quantum field theory.
@@
Let us now move on to general relativity. In the introduction to their paper, Chen & Goldstein, (2022) present the “Einstein equation (of general relativity)” (a.k.a. the Einstein field equations) as one of the motivations for adopting their account of laws, over a dynamic production account, because it is an equation that “in its usual presentation is non-dynamical.” Later, they return to the issue and explain that “There are ways of converting [the Einstein field equations] into FLOTEs [fundamental laws of temporal evolution] that are suitable for a dynamic productive interpretation.” Chen & Goldstein, (2022) give the ADM formalism as an example of a way this might be done (Arnowitt et al.,, 1962), complaining that this formalism discard solutions that are not globally hyperbolic (a bug that might be regarded as a feature if our universe is indeed globally hyperbolic and this formalism more precisely narrows down the space of physical possibilities [^47] See Maudlin, (2007, pg. 175, 189–191).). Along similar lines, Adlam, (2022a, sec. 2.2) writes

### v2

@@ 6.1 The evolution of the electromagnetic field
The wave equation for the scalar potential (17) fits the form of a three-dimensional wave equation (30) and thus we can formulate it (at the origin) as a spatiotemporally local law like (41),
@@
The wave equation for the vector potential (18) is three separate three-dimensional wave equations of the earlier form (30), one for each component of the potential. We can formulate it (at the origin) as
@@
For point charges, current density $\vec{J}$ that sources the vector potential can be formed by combining contributions of the form $\rho\vec{v}$ associated with each point charge, where $\rho$ is a delta function centered on the charge’s location and $\vec{v}$ is the charge’s velocity (Landau & Lifshitz,, 1971, sec. 28). As before, the velocity can be understood as a past derivative of the body’s location and thus depends on an arbitrarily short segment of its past trajectory. If the body has never moved faster than the speed of light, then this trajectory will lie within the past light-cone and the velocity will be a past light-cone neighborhood property. The current density will thus also be a past light-cone neighborhood property.
What if charged matter is modeled instead by a charge distribution? In that case, the current density at a moment cannot be read off from the present charge distribution and its recent history. Maudlin, (2018, pg. 8) gives the example of a uniform sphere that maintains a uniform charge density. This is compatible with there being no current density or with their being a current density corresponding to a rotation of the charge about some axis. Thus, unlike the velocity of a point charge, the current density of a continuum cannot be reduced to the historic behavior of charge. One option would be to treat the charge distribution and current density as distinct properties of the instantaneous state. The current density then becomes a space-time point property, not a past light-cone neighborhood property. Again, (43) counts as a spatiotemporally local dynamical law.
6.2 The forces on charged matter
@@
In words: the force from the electromagnetic field in a given direction depends on (i) the way the scalar potential is changing forward and backward in that direction along the past light-cone, (ii) the dot product of the charge’s velocity with the way the vector potential is changing forward and backward in that direction along the past light-cone, and (iii) the time derivative of the vector potential along the past trajectory of the charge. Here we see that (47) has the form of a non-dynamical law relating past light-cone neighborhood properties.
@@ 6.3 The reaction of matter to forces
Because the particles do not move faster than $c$, the momentum $\vec{p}^{\,p}$ will be a past light-cone neighborhood property and the right-hand side of (49), $\left(\frac{d}{dt}\right)^{f}\vec{p}^{\,p}$, will be a past-to-future light-cone neighborhood property. If we assume that we are dealing with purely electromagnetic forces then, as was discussed in section 6.2, the Lorentz force (44) that would appear on the right-hand side of (49) is a past light-cone neighborhood property. Thus, the reaction law (49) has the correct form to be a spatiotemporally local dynamical law. As in section 3, you could combine the non-dynamical Lorentz force law (47) and the dynamical reaction law (49) into a single dynamical law (assuming that the only forces at play are electromagnetic forces). Making that move, the laws of electromagnetism would all be dynamical laws and there would be no need for any non-dynamical laws.
@@
If the matter is modeled using continuous distributions of charge instead of point charges, the reaction of matter to forces is more complicated and will depend on the nature of the matter. You would need to couple some theory of the matter with the above laws of electromagnetism to get a complete theory that could fit the relativistic time evolution paradigm. A neatly relativistic way to do this would be to model the matter using a classical Klein-Gordon or Dirac field in a Maxwell-Klein-Gordon or Maxwell-Dirac classical field theory. Then, one would need to show that the wave equations for these fields can be formulated in ways that fit the relativistic time evolution paradigm, as we have seen that the wave equations of electromagnetism can—(42) and (43). For the Klein-Gordon field, that is straightforward as we can simply carry over the techniques that were used for the electromagnetic field. [^39] You can prove the same kind of causality theorem for the Klein-Gordon equation as the one we discussed for the wave equation near the beginning of this section (Wald,, 1984, sec. 10.1; Strauss,, 2008, pg. 234, problem 8). For the Dirac field, it is not so clear how to formulate the wave equation as a spatiotemporally local dynamical law. Note that with either the Klein-Gordon or Dirac equation in place, the Lorentz force law (48) is not needed as a fundamental non-dynamical law (though it could still be regarded as a non-fundamental law giving the density of force exerted by the electromagnetic field on the Klein-Gordon or Dirac field [^40] The idea that fields can experience forces is discussed in Sebens, (2018, 2021).).
6.4 The whole package
@@
Putting it all together, we have arrived at a way of formulating electromagnetism that seems like it could fit the relativistic time evolution paradigm. For point charges, we can take the laws to be given by the wave equations for the potentials, (42) and (43), the Lorentz force law, (47), and the relativistic version of Newton’s second law (49)—or, (47) and (49) could be combined. However, there is danger on the horizon for that kind of theory. We could run into problems of self-interaction like those faced by Newtonian gravity with point masses (discussed in section 3).
Just as the gravitational field becomes infinite as you approach a point mass and ill-defined at its location, the electric field becomes infinite as you approach a point charge and ill-defined at its location. This causes trouble for the standard Lorentz force law (14). Does our new formulation of the force law (47) face the same problems? For a single point charge $q$ at rest, we can use the scalar potential $\phi=\frac{q}{r}$ and set the vector potential to zero. The force law (47) yields
@@
We have seen that this formulation of electromagnetism with point charges satisfies the first two conditions of the relativistic time evolution paradigm: the dynamical and non-dynamical laws take the correct form. We have not yet settled whether the formulation satisfies the third condition as we have not yet settled whether the theory is locally deterministic. To do so, we must figure out whether, in general, an arbitrarily thin law-abiding time slice of a contracting light-cone yields a unique remainder of that light-cone. This can be proven for the wave equations governing the evolution of the fields, [^43] See the references in footnote 31. but proving it for the entire theory requires looking at the coupled equations for the field and particle dynamics. [^44] See Frisch, (2004, sec. 2); Frisch, (2005, pg. 32–35). Doing this honestly would require a solution to the aforementioned problems of self-interaction. There is another complication as well. Hartenstein & Hubert, (2021) have shown that specifying law-abiding combinations of electromagnetic field and point charge states at an instant generically yields infinities or discontinuities that propagate along the future light-cone (called “shock fronts”) and lead to ill-defined dynamics. [^45] See also Lazarovici, (2018, sec. 8.1). These shock fronts occur because the arbitrarily chosen field around a point charge does not capture the way that the charge’s past motion would have acted as a source for the field. For example, at some moment you might have a particle with an initial velocity surrounded by the Coulomb field of a stationary point charge. That is not compatible with any reasonable past. Our formulation of determinism (from section 2) requires that we start with an arbitrarily short time interval where the laws are obeyed and that starting point should help in addressing these problems. We are assuming that the interaction between particle and field has been well-behaved in the immediate past and then asking whether the future evolution is well-behaved and unique.
@@
Shifting now to consider electromagnetism with a continuous charge distribution, the wave equations for the potentials, (42) and (43), would again be spatiotemporally local dynamical laws. In addition, we would need law(s) determining the evolution of the charged matter. The Lorentz force density law (48) can be formulated as an acceptable non-dynamical law, but it will not be a fundamental law for certain classical completions of electromagnetism. In particular, if we model the evolution of matter using the Klein-Gordon or Dirac equations in Maxwell-Klein-Gordon or Maxwell-Dirac field theory, it is not needed. As was discussed earlier, the Klein-Gordon equation can be formulated as a spatiotemporally local dynamical law and one might find a way to do so for the Dirac equation. The Maxwell-Klein-Gordon and Maxwell-Dirac field theories could then satisfy the first two conditions of the relativistic time evolution paradigm. They should satisfy the third condition as well—the free Klein-Gordon and Dirac equations are locally deterministic [^46] See Wald, (1984, sec. 10.1); Thaller, (1992, sec. 1.5); Strauss, (2008, pg. 234, problem 8) and the introduction of interactions does not bring about any infinite or ill-defined self-interaction. [^47] See Sebens, (2023).
@@ 7 The Path Ahead
We have thus far focused our attention on parts of physics that seem to fit well with the dynamic production account of laws (Newtonian gravity and electromagnetism), noting the challenges that arise there and exploring how those challenged might be addressed. Let us now turn briefly to two physical theories that are more hostile to an interpretation in terms of dynamic production: quantum field theory and general relativity. The purpose of this quick treatment is just to suggest that the dynamic production account has a shot at being extended to our most successful physical theories, and should not be viewed as roadblocked in a way that would make the kind of project pursued here a futile exercise (a worry that was raised at the end of section 2).
@@
To build up to quantum field theory, let us start by discussing the non-relativistic quantum mechanics of a fixed number of particles. In that context, there exist a handful of “interpretations of quantum mechanics” that give explicit proposals about the laws and ontology—such as GRW theory, Bohmian mechanics, the many-worlds interpretation, and the many interacting worlds approach. These four options can all be fit straightforwardly into the time evolution paradigm, though there are other “retrocausal,” “all-at-once,” or “non-Markovian” options [^48] See Wharton, (2014); Adlam, (2018, 2022a, 2022b, 2023); Friederich & Evans, (2023); Barandes, (2023). that cannot. Focusing on the four aforementioned options, the dynamics will be either first-order or second-order depending on the interpretation. The many-worlds interpretation simply takes the first-order Schrödinger equation to give the dynamics of the wave function. This is naturally interpreted as taking the wave function at a moment as input (with no need to look to the arbitrarily-short past) and returning as output the way the wave function changes into the arbitrarily short future (Easwaran,, 2014, pg. 857). GRW theory introduces stochastic exceptions to the Schrödinger equation that can be understood as giving probabilities for different behavior in the arbitrarily short future (Sebens,, 2015a, sec. 2). Bohmian mechanics is often presented as adding a first-order guidance equation, that is another law beyond the Schrödinger equation, determining the velocities of particles from their locations and the wave function (velocities that could be interpreted as future neighborhood properties). However, a second-order guidance equation can be used instead provided we impose the first-order guidance equation as a restriction on initial conditions (Goldstein & Struyve,, 2015; Dewdney,, 2023). In the many interacting worlds approach, the dynamics is second-order as particles are reacting to forces by Newton’s second law and (at the fundamental level) there is no wave function evolving by the Schrödinger equation (Hall et al.,, 2014; Sebens,, 2015b; Ghadimi et al.,, 2018). The fact that quantum theories normally posit first-order dynamics has been used by Builes & Impagnatiello, (forthcoming) to argue that we live in a world where the present moment is enough to produce the future (the arbitrarily-short past is not needed as input)—in their words, “our universe is Markovian.” I prefer to leave open whether the fundamental dynamics will turn out to be first-order or second-order.
@@
Deciding between these competing interpretations of quantum mechanics is difficult. The many-worlds interpretation has been touted as “the only game in town” because it is the only option that can be immediately extended to relativistic quantum field theory (Wallace,, 2012, pg. 35; Wallace,, 2023). That virtue could be challenged either because the many-worlds interpretation has other serious problems (Adlam,, 2023) or because it is not so easy to formulate quantum field theory in terms of a quantum state evolving by a Schrödinger equation (Sebens,, 2022a). [^49] Adlam, (2022a, sec. 2.1) takes the use of path integrals in quantum field theory to suggest that the theory might best be understood outside the time evolution paradigm. Although it may be possible to use path integrals to formulate the theory in a way that violates the time evolution paradigm, path integrals in quantum field theory can be viewed as tools for calculating state evolution. The problem for Schrödinger evolution is that it is unclear whether the quantum state is a particle wave function (in Fock space) or a field wave functional, though hopefully one of these proposals can be made to work. Although the other interpretations of quantum mechanics may not be as straightforward to extend to quantum field theory as the many-worlds interpretation, there exist a number of promising strategies for doing so. The attempts to extend GRW and Bohmian mechanics to quantum field theory include non-local interactions that are incompatible with the relativistic time evolution paradigm (Maudlin,, 2019, ch. 7). This is to be expected, as it is the lesson of EPR and Bell’s theorem (taken along with the relevant experimental tests and making certain natural assumptions) that any single-world quantum theory must include such non-local interactions (Maudlin,, 2014, 2019). One might react to this situation by abandoning the relativistic time evolution paradigm and instead simply seeking theories that fit the original time evolution paradigm, accepting that dynamic production occurs relative to some preferred foliation. Alternatively, one might work to show that a version of quantum field theory with multiple worlds can be cast in a form that fits the relativistic time evolution paradigm (adopting the many-worlds interpretation or the many interacting worlds approach). Of course, one could also react to this situation by abandoning dynamic production (Adlam,, 2023). I simply want to highlight that there are ways to potentially retain dynamic production in quantum field theory.
@@
Let us now move on to general relativity. In the introduction to their paper, Chen & Goldstein, (2022) present the “Einstein equation (of general relativity)” (a.k.a. the Einstein field equations) as one of the motivations for adopting their account of laws, over a dynamic production account, because it is an equation that “in its usual presentation is non-dynamical.” Later, they return to the issue and explain that: “There are ways of converting [the Einstein field equations] into FLOTEs [fundamental laws of temporal evolution] that are suitable for a dynamic productive interpretation.” Chen & Goldstein, (2022) give the ADM formalism as an example of a way this might be done (Arnowitt et al.,, 1962), complaining that this formalism discards solutions that are not globally hyperbolic (a bug that might be regarded as a feature if our universe is indeed globally hyperbolic and this formalism more precisely narrows down the space of physical possibilities [^50] See Maudlin, (2007, pg. 175, 189–191).). Along similar lines, Adlam, (2022a, sec. 2.2) writes